# Dataset Distillation as Pushforward Optimal Quantization

### v2

@@ 7 $i\leftarrow i+1$;
Convergence to an optimal quantizer is only guaranteed in the case of a log-concave distribution in one dimension [21, 27]. However, in higher dimensions, convergence to a (not necessarily optimal) stationary grid is possible [35, 34].
2.3 Score-based diffusion
To connect the quantization error on the latent space with the quantization error in the image space, we need to consider properties of the latent-to-image process. In particular, we focus on score-based diffusion models, interpreted as discretizations of particular noising SDEs [49]. Consider the following SDE, where $W$ is a standard Wiener process:
[…]
Then, where the density of $x$ at time $t>0$ is given by $p_{t}$, the reverse of the diffusion process is given by the reverse-time SDE [16],
[…]
where $\bar{W}$ is the standard Wiener process running in reverse time from time $T$ to 0. For an increasing noising schedule $\sigma(t)\in[0,+\infty)$ or noise-scale $\beta(t)\in[0,1)$, the variance-exploding SDE (VESDE, or Brownian motion) is given by
[…]
and the variance-preserving SDE (VPSDE, or Ornstein–Uhlenbeck process) is given by
[…]
The VPSDE and VESDE are commonly used when training diffusion models due to their simple discretization. Indeed, they both arise as continuous limits of earlier proposed diffusion models. Specifically, VPSDE corresponds to denoising score matching with Langevin dynamics [48], and VESDE to denoising diffusion probabilistic models [47, 17].
Suppose that the true data distribution on the image space is given by $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with bounded support. [39] demonstrates that the backward SDE processes are indeed diffusions for the VPSDE and VESDE. Moreover, the diffusion satisfies a data-processing inequality, showing that the $f$-divergence after backwards diffusion is at most the $f$-divergence at marginal time $T$. We review this in Appendix˜D.
Convergence in $f$-divergence necessarily requires absolute continuity of the compared marginal with respect to the underlying diffused distribution, which is equivalent to continuity with respect to the Lebesgue measure by the Hörmander condition. This rules out singular initializations such as empirical measures, as one might want to use for dataset distillation. To combat this, we use weak convergence with appropriately chosen test functions.
The goal in question: given Wasserstein-2 convergence of the marginals $\nu_{T}^{(k)}\rightarrow\mu_{T}$, we wish to derive a bound on expectations $\mathbb{E}_{\nu_{\delta}^{(k)}}[f]\xrightarrow{?}\mathbb{E}_{\mu_{\delta}}[f]$, for some fixed $\delta\in(0,T)$ and $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ satisfying some regularity conditions. Such a bound would directly link to training neural networks with surrogate data, e.g. by taking $f$ to be the gradient of a loss function with respect to some network parameters.
@@ 2.3 Score-based diffusion
Having $\delta=0$ may not be well defined because of non-smoothness and blowup of the score at time 0 for singular measures [39, 54].
@@ 3 Dataset Distillation as Optimal Quantization
Our main result Theorem˜1 gives consistency of dataset distillation for a score-based diffusion prior in the image space. It transforms convergence of the terminal distribution into convergence of the backwards-diffused distributions at small time. Later, we use this in Section˜3.1 to present DDOQ as a modification of the D4M method. By simply changing the clustering objective from a Wasserstein barycenter to an optimal quantization by adding weights, we can effectively reduce the Wasserstein distance to the data distribution.
@@
Consider the VESDE/Brownian motion
[…]
or the VPSDE/Ornstein–Uhlenbeck process
@@
Theorem˜1 combined with the asymptotic rates of [13] gives convergence rates as the number of quantization points increases. We note that this can be further be combined with the convergence of optimal quantization rates of empirical measures such as Theorem˜3 in Section˜C.2.
@@
In particular, as the number of points increases, we have convergence to the underlying data distribution in image space, giving consistency.
@@ 3.1 Proposed method: Dataset Distillation by Optimal Quantization
We have seen that clustering gives a consistent approximation to the latent distribution. Moreover, using the generative diffusion model, this then gives a consistent approximation in the image space. Now we present the proposed algorithm, Dataset Distillation by Optimal Quantization (DDOQ). We detail the steps in text below and summarize in Algorithm˜2.
Suppose we are given an LDM, i.e. a (conditional) encoder-decoder pair and a diffusion-based model on the latent space [43]. Let $K$ be the target number of images per class. Constructing the distilled dataset consists of the following steps in sequence.
@@ 4 Experiments
We compare with two different latent diffusion architectures, namely the original latent diffusion model utilizing UNets [43], then the stronger diffusion transformer (DiT) architecture [38]. We first use the LDM to compare the pure algorithmic differences of DDOQ with D4M and related baselines. Then, we use DiT to demonstrate the potential of DDOQ compared to newer state-of-the-art baselines.
@@ 4.1 UNet backbone
Table 1: Comparison of top-1 classification performance on the ImageNet-1K dataset for baselines versus the proposed DDOQ method at various IPCs. We observe that DDOQ outperforms the previous SOTA method D4M, due to the addition of weights to the synthetic data. The maximum performance for all methods should be 69.8 as the soft labels are computed using a pre-trained ResNet-18 model. Figures for TESLA, SRe2L and D4M are taken directly from [50], RDED and CDA from [51, 55].
@@
To validate the proposed DDOQ algorithm, we directly compare with the previous state-of-the-art methods D4M [50] and RDED [51] on the ImageNet-1K dataset. The latent diffusion model chosen for latent generation and image synthesis is the publicly available pre-trained Stable Diffusion V1.5 model. We report results for SRe2L [56], RDED [51], CDA [55] as given. For low IPC, we also report the TESLA method, which is a SOTA bi-level method based on MTT [7]. RDED achieves SOTA results for low IPC using its aggregated images and special training schedule, while CDA improves upon SRe2L using time-varying augmentation.
For consistency and a more direct comparison with previous methods, we use the pre-trained PyTorch ResNet-18 model to compute the soft labels, using the same protocol as [50]. After computing the soft labels using the pre-trained ResNet-18 model, we train new ResNet-18, ResNet-50 and ResNet-101 models to match the soft labels. The data augmentation is also identical, with the only difference being the addition of the weights to the training objective and minimal optimizer hyperparameter tuning for the new objective. We provide a direct comparison of the distilled data performance in Table˜1 for the IPCs $K\in\{10,50,100,200\}$. Our figures for DDOQ are averaged over five models trained on the same distilled data.
@@
Table 2: Generalization performance of D4M and the proposed DDOQ method for different soft-label teachers and student architectures at IPC 50. We observe that DDOQ has uniformly better cross-architecture generalization for convolutional teacher and student architectures, and slightly worse performance for student models using the transformer architecture Swin-T. Figures for D4M as reported in [50]. Bold indicates best method.
@@
Table 3: Wasserstein-2 distances of the distilled latents compared to the encoded training data. We observe the weighting drops the Wasserstein-2 distance by around 20%.
Test class 0 1 2
IPC 10 D4M 49.67 51.14 47.16
DDOQ 42.97 42.71 39.14
IPC 50 D4M 47.34 49.10 45.28
DDOQ 40.41 41.01 37.53
Table˜2 shows the generalization performance of the distilled latent points. Different PyTorch pre-trained teacher networks are used to create the soft labels as in Step 3 of DDOQ, in particular MobileNet-V2 and Swin-T. We then evaluate the distilled images and soft labels with three convolutional architectures and the Swin-T transformer architecture. We observe that DDOQ is not only able to generalize to different model architectures, but also uniformly outperforms D4M on all convolutional student architectures. The slightly worse performance of DDOQ when using a transformer architecture for the student may be due to more precise hyperparameter tuning requirements.
To illustrate the weights, Figure˜1 plots ten example images from the “jeep” class when distilled using $K=10$ IPC. We observe that there is a very large variance in the weights $v_{k}^{(L)}\big{/}\sum_{j=1}^{K}v_{j}^{(L)}$}, indicating the presence of strong clustering in the latent space. Nonetheless, there is no qualitative evidence that the weights indicate “better or worse” training examples. Indeed, among other classes, there are synthetic images that do not lie in the appropriate class, but are still weighted highly.
To consider the effect of weighting the centroids, we consider the class-wise Wasserstein distance between the centroids and encoded training data, with and without the weights. Table˜3 plots the Wasserstein distance between the distilled latent points and the encoded training dataset, for the first three classes of ImageNet-1K. The latent points are clustered with uniform weights (D4M), and with variable weights (DDOQ). We observe that adding weights reduces the Wasserstein distance uniformly by around 20%. This leads to better gradient faithfulness when training a neural network.
@@ 4.2 DiT backbone
To compare the potential of DDOQ, we use a stronger generative model, namely the diffusion transformer [38]. This architecture achieves uniformly better sample quality compared to the LDM used in the previous section, in terms of Inception Score and Fréchet inception distance.
We compare with the SOTA DD methods based on diffusion guidance, namely Minimax [14] and Influence Guided Diffusion (IGD) [6], which both use DiT. These methods guide the decoding process using some fine-tuning or batch statistics. In contrast, DDOQ directly modifies the initialization in the latent space. In addition, we compare with samples directly generated with DiT from random initializations. We compare on ImageNet-1K, as well as the 10-class subsets ImageNette and ImageWoof. The baseline figures are taken from [6] which report higher numbers than [14].
Table˜4 demonstrates that we significantly outperform the diffusion-guidance based methods on the full ImageNet-1K, as well as DDOQ with the UNet backbone. Moreover, we are competitive-with or better than the baselines on the 10-class subsets, namely outperforming in the low IPC setting and in the more difficult ImageWoof subset.
Table 4: Comparisons on ImageWoof and ImageNette using the ResNetAP-10 architecture, and ImageNet-1K using ResNet-18. We observe that DDOQ-DiT is able to outperform the guided diffusion methods at low IPC and are competitive at higher IPC, namely outperforming the compared methods on ImageNet-1K. The maximum accuracies are 87.2, 94.6 and 69.8 for ImageWoof, ImageNette and ImageNet-1K respectively.
@@ 5 Conclusion
This work proposes DDOQ, a dataset distillation method based on optimal quantization. Inspired by optimal quantization and Wasserstein theory, we theoretically demonstrate consistency of the distilled datasets in Theorem˜1 when using standard diffusion-based generative models to generate the synthetic data. We empirically demonstrate performance higher than previous state-of-the-art methods on a large-scale ImageNet-1K dataset, as well as in the low IPC setting of ImageNet subsets.
We have presented theoretical justification for one framework of dataset distillation. Future work could include sharper bounds in Theorem˜1 that exploit the sub-Gaussianity of the diffused distributions or manifold hypothesis, or investigating alternative choices of weight variance reduction. Other interesting directions could be relating the weightings of the synthetic data to the hardness of learning the data, such as in [20], further increasing the performance using different training regimes such as curriculum learning, or extending the theoretical analysis to inexact score matching. Possible empirical work could investigate correlations between the weights and influence [42].
Acknowledgments and Disclosure of Funding
@@ Appendix A Details on heuristic bi-level distillation
    (Meta-learning.) This uses the assumption that more data produces better results, replacing the risk matching objective with a risk minimization objective and removing the dependence on $\mathcal{T}$ [53, 10]. The learning algorithm $\Phi_{S}=\Phi_{S}(\theta_{0})$ with initial parameter $\theta_{0}$ is given by unrolling gradient steps $\theta_{t+1}=\theta_{t}-\eta\nabla\mathcal{L}_{S}(\theta_{t})$. The distilled dataset minimizes the training loss as trained on the distilled set up to $T$ iterations:
@@
    (Distribution matching.) Inspired by reproducing kernel Hilbert spaces and the empirical approximations using random neural network features, [58] proposes an alternate distillation objective that is independent of target loss. The minimization objective is
@@
    (Trajectory matching.) For a fixed network architecture, this method aims to match the gradient information of the synthetic and true training datasets from different initializations [5]. The heuristic is that similar network parameters will give similar performance. In addition, the gradients are allowed to be accelerated by matching a small number of synthetic training steps with a large number of real data training steps: for $N\ll M$ steps, the Matching Training Trajectory (MTT) objective is (with abuse of notation):
@@
Appendix B Proof of Proposition 1
@@
Suppose we have a quantization ${\mathbf{x}}=\{x_{1},...,x_{K}\}$. Assume that the (probability) measure $\mu$ is null on the boundaries of the Voronoi clusters $\mu(\partial(C_{i}))=0$. Then the measure $\nu$ that minimizes the Wasserstein-2 distance 2 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$. Moreover, the optimal coupling is given by the projection onto the centroids.
@@
where the first inequality comes from definition of distance to a set ($\gamma$-a.s.) and the support condition on $\nu$, and the equality from the marginal property of couplings. The final term is attained when $\nu$ has the prescribed form: the coupling is $c(x_{i},y)=\mathbf{1}_{y\in C_{i}}$ and the corresponding transport map is projection onto the set $\mathbf{x}$} (defined $\mu$-a.s. from the null condition). This shows a lower bound of 2 that is attained. ∎
Appendix C Optimal quantization
C.1 Proof of Thm 2

### v3

@@ 2.2 Score-based diffusion
Having $\delta=0$ may not be well defined because of non-smoothness and blowup of the score at time 0 for singular measures (Pidstrigach, 2022; Yang et al., 2023).
@@
Working with weak convergence is necessary due to the singular empirical measures. Pidstrigach (2022) demonstrates that the backward SDE process satisfies a data-processing inequality, showing that the $f$-divergence after backwards diffusion is at most the $f$-divergence at marginal time $T$. However, $f$-divergences require absolute continuity of the compared marginal with respect to the underlying diffused distribution, which is equivalent to absolute continuity with respect to the Lebesgue measure by the Hörmander condition. This rules out singular initializations such as empirical measures, which arise in dataset distillation.
@@ 3 Dataset Distillation as Optimal Quantization
Our main result Theorem˜1 gives consistency of dataset distillation for a score-based diffusion prior in the image space. Later, we use this in Section˜3.1 to present DDOQ as a modification of the D4M method. By simply changing the clustering objective from a Wasserstein barycenter to an optimal quantization by adding weights, we can effectively reduce the Wasserstein distance to the data distribution. Moreover, from Proposition˜3, the weights are automatically determined during the $k$-means clustering process.
@@
Consider the VESDE/Brownian motion or the VPSDE/Ornstein–Uhlenbeck process
@@
Theorem˜1 combined with the asymptotic rates of Graf and Luschgy (2000) gives convergence rates as the number of quantization points increases. We note that this can be further be combined with the convergence of optimal quantization rates of empirical measures such as Theorem˜3 in Section˜C.4. In particular, the next result shows that as the number of points increases, we have convergence to the underlying data distribution in image space, giving consistency.
@@ 6 To train a new network $f_{\theta}$:
Table 1: Wasserstein-2 distances of the distilled latents compared to the encoded training data on classes of ImageNet-1K. We observe the weighting drops the Wasserstein-2 distance significantly.
Test class 0 1 2 Avg reduction
IPC 10 D4M 49.67 51.14 47.16 -15.7%
DDOQ 42.97 42.71 39.14
IPC 50 D4M 47.34 49.10 45.28 -16.1%
DDOQ 40.41 41.01 37.53
@@ 3.1 Proposed method: Dataset Distillation by Optimal Quantization
We have seen that clustering gives a consistent approximation to the latent distribution. Using the generative diffusion model and Theorem˜1, we obtain a consistent approximation to the original data distribution in the image space. We now propose Dataset Distillation by Optimal Quantization (DDOQ). We detail the steps in text below and summarize in Algorithm˜1.
Suppose we are given a latent diffusion model (LDM), i.e. a (conditional) encoder-decoder pair and a diffusion model on the latent space (Rombach et al., 2022). Let $K$ be the target number of images per class. Constructing and using the distilled dataset consists of the following steps in sequence.
@@
Within Figure˜1, the encoder and decoder are given by a pre-trained encoder-decoder model, combined with a diffusion model in the latent space. Compared with D4M, we include (automatically determined) weights when training a new network in the final step, indicated by the variable length arrows. These are justified by considering the expectation of network gradients with respect to the training and distilled distributions. As seen in Table˜1, the inclusion of the weights significantly decreases the Wasserstein distance of the distilled points to the data distribution, when tested on classes of ImageNet-1K.
@@ 4 Experiments
We compare with two different latent diffusion architectures, namely the original latent diffusion model utilizing UNets (Rombach et al., 2022), then the stronger diffusion transformer (DiT) architecture (Peebles and Xie, 2023). We first use the LDM to compare the pure algorithmic differences of DDOQ with D4M and related baselines. Then, we use DiT to demonstrate the potential of DDOQ compared to newer state-of-the-art baselines.
@@ 4.1 UNet backbone
Table 2: Comparison of top-1 classification performance on the ImageNet-1K dataset at various IPCs. We observe that the proposed DDOQ outperforms the clustering-based method D4M, due to the addition of weights to the synthetic data. The maximum performance for all methods should be 69.8 as the soft labels are computed using a pre-trained ResNet-18 model. In particular, we achieve a 30% reduction in error gap using ResNet-101 at IPC 200.
@@
To validate the proposed DDOQ algorithm, we directly compare with the previous state-of-the-art disentangled methods D4M (Su et al., 2024) and RDED (Sun et al., 2024) on the ImageNet-1K dataset. Baseline figures are reported as given in their respective works. For low IPC, we also report the TESLA method, which is a SOTA bi-level method based on MTT (Cui et al., 2023), but is unscalable past IPC=10. RDED achieves strong results for low IPC using its aggregated images and special training schedule, while CDA improves upon SRe2L using time-varying augmentation. For consistency, soft labels are computed using the pre-trained PyTorch ResNet-18 model, and new ResNet-{18,50,101} models are trained using the distilled data. We provide a direct comparison of the distilled data performance in Table˜2 for the IPCs $K\in\{10,50,100,200\}$. Variances for DDOQ are averaged over five models trained on the same distilled data.
@@
Table 3: Generalization performance of D4M and the proposed DDOQ method for different soft-label teachers and student architectures at IPC 50. We observe that DDOQ has uniformly better cross-architecture generalization for convolutional teacher and student architectures, and slightly worse performance for student models using the transformer architecture Swin-T.
@@
Table˜3 shows the generalization performance of the distilled latent points. The PyTorch pre-trained MobileNet-V2 and Swin-T networks are used to create the soft labels. We then evaluate the distilled images and soft labels with three convolutional architectures and the Swin-T transformer architecture. We observe that DDOQ is not only able to generalize to different model architectures, but also uniformly outperforms D4M on all convolutional student architectures. The slightly worse performance of DDOQ when using a transformer architecture for the student may be due to more precise hyperparameter tuning requirements.
To illustrate the weights, Figure˜2 plots ten example images from the “jeep” class when distilled using $K=10$ IPC. We observe that there is a very large variance in the weights $v_{k}^{(L)}\big/\sum_{j=1}^{K}v_{j}^{(L)}$}, indicating the presence of strong clustering in the latent space. Nonetheless, there is no qualitative evidence that the weights indicate “better or worse” training examples, rather indicating the structure in the latent space.
@@ 4.2 DiT backbone
To compare the potential of DDOQ, we use a stronger generative model, namely the diffusion transformer (DiT) (Peebles and Xie, 2023). This architecture achieves uniformly better sample quality compared to the LDM used in the previous section, in terms of Inception Score and Fréchet inception distance. We denote the method with DiT backbone as DDOQ-DiT.
We compare with the SOTA DD methods based on diffusion guidance, namely Minimax (Gu et al., 2024) and Influence Guided Diffusion (IGD) (Chen et al., 2025), which both use DiT. These methods guide the decoding process using some fine-tuning or batch statistics. In contrast, DDOQ-DiT directly modifies the initialization in the latent space. In addition, we compare with samples directly generated with DiT from random initializations, labelled ‘DiT’. We compare on ImageNet-1K, as well as the 10-class subsets ImageNette and ImageWoof. The baseline figures are taken from Chen et al. (2025) which report higher numbers than Gu et al. (2024).
Table˜4 demonstrates that we significantly outperform the diffusion-guidance based methods on the full ImageNet-1K, as well as DDOQ-DiT with the UNet backbone. Moreover, we are competitive-with or better than the baselines on the 10-class subsets, namely outperforming in the low IPC setting and in the more difficult ImageWoof subset. Moreover, the stronger diffusion model significantly benefits the low IPC ImageNet-1K setting, increasing ResNet-18 test accuracy from 33.1 to 53.0. An ablation on the effect of different weights during training is given in Appendix˜I.
Table 4: Comparisons on ImageWoof and ImageNette using the ResNetAP-10 architecture, and ImageNet-1K using ResNet-18. DDOQ-DiT outperforms the guided diffusion methods at low IPC and is competitive at higher IPC, namely outperforming the compared methods on ImageNet-1K.
@@ 5 Conclusion
This work proposes DDOQ, a dataset distillation method based on optimal quantization. Inspired by optimal quantization and Wasserstein theory, we theoretically demonstrate consistency of the distilled datasets in Theorem˜1 when using standard diffusion-based generative models to generate the synthetic data. Experiments show the proposed method is competitive or better than SOTA on large-scale ImageNet experiments.
We have presented theoretical justification for disentangled dataset distillation methods, which rely on clustering as the mechanism for approximating the underlying data distributions. More specifically, we justify the combination of a low-dimensional latent space and a diffusion model. Consistency or convergence of other dataset distillation frameworks such as bi-level methods and diffusion guidance are still open questions.
Future work could include sharper bounds in Theorem˜1 that exploit the sub-Gaussianity of the diffused distributions or manifold hypothesis. Other interesting directions could be relating the weightings of the synthetic data to the hardness of learning the data, such as in (Joshi and Mirzasoleiman, 2023), further increasing the performance using different training regimes such as curriculum learning, or extending the theoretical analysis to inexact score matching. Possible empirical work could investigate correlations between the weights and influence (Pruthi et al., 2020).
Acknowledgments
@@ Appendix A Details on heuristic bi-level distillation
    (Meta-learning.) This uses the assumption that more data produces better results, replacing the risk matching objective with a risk minimization objective and removing the dependence on $\mathcal{T}$ (Wang et al., 2018; Deng and Russakovsky, 2022). The learning algorithm $\Phi_{S}=\Phi_{S}(\theta_{0})$ with initial parameter $\theta_{0}$ is given by unrolling gradient steps $\theta_{t+1}=\theta_{t}-\eta\nabla\mathcal{L}_{S}(\theta_{t})$. The distilled dataset minimizes the training loss as trained on the distilled set up to $T$ iterations:
@@
    (Distribution matching.) Inspired by reproducing kernel Hilbert spaces and the empirical approximations using random neural network features, Zhao and Bilen (2023) proposes an alternate distillation objective that is independent of target loss. The minimization objective is
@@
    (Trajectory matching.) For a fixed network architecture, this method aims to match the gradient information of the synthetic and true training datasets from different initializations (Cazenavette et al., 2022). The heuristic is that similar network parameters will give similar performance. In addition, the gradients are allowed to be accelerated by matching a small number of synthetic training steps with a large number of real data training steps: for $N\ll M$ steps, the Matching Training Trajectory (MTT) objective is (with abuse of notation):
@@
Appendix B Additional definitions
The Wasserstein-2 distance is defined as
Appendix C Optimal quantization
C.1 Additional background
Optimal quantizers converge in Wasserstein distance at a rate $\Theta(K^{-1/d})$ to their respective distributions as the number of particles $K$ increases (Graf and Luschgy, 2000). Moreover, any limit point of the optimal quantizers is an optimal quantizer of the limiting distribution (Pollard, 1982). Moreover, it can be shown that if a sequence of distributions converges in Wasserstein distance $\mu_{n}\rightarrow\mu$, then so do the errors in quantization for a fixed quantization level (Liu and Pagès, 2020, Thm. 4).
C.2 Proof of Proposition 1
@@
Suppose we have a quantization ${\mathbf{x}}=\{x_{1},...,x_{K}\}$. Assume that the (probability) measure $\mu$ is null on the boundaries of the Voronoi clusters $\mu(\partial(C_{i}))=0$. Then the measure $\nu$ that minimizes the Wasserstein-2 distance 14 and satisfies $\operatorname{supp}\nu\subset\{x_{1},...,x_{K}\}$ is $\nu_{K}=\sum_{i=1}^{K}\mu(C_{i})\delta(x_{i})$. Moreover, the optimal coupling is given by the projection onto the centroids.
@@
where the first inequality comes from definition of distance to a set ($\gamma$-a.s.) and the support condition on $\nu$, and the equality from the marginal property of couplings. The final term is attained when $\nu$ has the prescribed form: the coupling is $c(x_{i},y)=\mathbf{1}_{y\in C_{i}}$ and the corresponding transport map is projection onto the set $\mathbf{x}$} (defined $\mu$-a.s. from the null condition). This shows a lower bound of 14 that is attained. ∎
C.3 Proof of Proposition 2